Three individuals form a partnership and agree to divide the profits equally. X invests
$9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less
does x receive compared to having the profits divided in proportion to the amounts
invested by X, Y, and Z?

Answer by rfer(16322)

9x+7x+4x=4800
20x=4800
x=240
-----
X is 9*240=2160
y is 7*240=1680
Z is 4*240=960
-------------
4800/3=1600 equal amounts